Calculate the iterated integral. $$***/x2 + yż dy dx 2 (2017) – ) - 150 Need Help? Read It
Answer #1

w * A fls zey I waya dulux A = dy da -O where x as constant Put x² + y² = t • Diff. w.r.t.y 0+ 2y =dt ydy= when y=o t= x²Now put x²+1 = u 2 uto= du (x=0 - u=1 42 = 37uaw) ada du da lo Ź * ((10) 512-) = Į ( 105h-1) 243 da - s A = (5 (1954-)) - 3 (